A car moving with a velocity of 54 km/h accelerate uniformly at the rate of 2 m/s². Calculate the time taken to cover this distance

1
Expert's answer
2022-05-26T14:07:10-0400
"a=\frac{\Delta v}{\Delta t}"

"\Delta t=\frac{\Delta v}{a}=\frac{(72-54)\div 3.6\:\rm m/s}{2\:\rm m/s^2}=2.5\:\rm s"